OpenStudy (anonymous):

olve the following system of equations. Express your answer as an ordered pair in the format (a,b), with no spaces between the numbers or symbols. 2x+7y=-1 4x-3y=-19

OpenStudy (anonymous):

I would probably use elimination to solve this one. 2x+7y=−1 4x−3y=−19 Multiply the first equation by -2 and add it to the 2nd equation. −2(2x)+4x−2(7y)−3y=−2(−1)−19 −17y=−17 Now solve for y. Then plug in your answer for y into either the 1st or 2nd equation to find x.

OpenStudy (anonymous):

ok thank you